Grade 12th passPhysical Chemistry

hello, could you please help me with this?
butene gas readily burns with oxygen according to the following equation C4H8 + 6O2 --> 4CO2 + 4H2O
What volume of butene at 188 ⁰C and 2.50 atm can be burned with 12.0 L of O2 at 745 torr and 25.0 ⁰C?

ninigooz
10 Years agoGrade 12th pass

1 Answer

Ashutosh Mohan Sharma
10 Years ago
moles of O2 = PV/RT = [(745/720)* 12]/[0.082 * (25+273)] = 12.2769
moles of butene = 1/6 moles of O2
therefore moles of butene =2.0461
2.0461= 2.5*V / 0.082* (188+273)
solve to get value of V .